TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-04-25 23:02:00
subject: 4\11 Pt 1 ISS On Orbit Status 11-04-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

12 Apr 2003

ISS On-Orbit Status 11 Apr 2003

Part 1 of 2

All ISS systems continue to function nominally, except as noted
previously or below.  Space Day 139 for Expedition 6 (since STS-113
launch on 11/23/02).

FE-1 Nikolai Budarin conducted another session with the Russian
MedOps cardio experiment MO-1 (Study of the Bioelectric Activity of
the Heart at Rest), with Don Pettit assisting as CMO (crew medical
officer). [During the test, the data of the Gamma-1M ECG
(electrocardiogram) were recorded for later downlink.]

Budarin also performed more preventive maintenance in the FGB, today
replacing its ten smoke detectors (IDZ-2), which have reached the end
of their certified lifetime.  They are subunits of the FGB fire
detection and suppression system (SPOPT).  [The FGB has nine IDZs in
the instrumentation cargo compartment (PGO) and one in the
pressurized adapter (GA), docked to PMA-1 (pressurized mating adapter
1) and the Node.  These 184 x 130 x 102 mm units work by first
ionizing air molecules inside their charge chamber with a small
amount of radioactive material and then measuring the conductivity of
the air.  Air conductivity is reduced if it contains smoke particles,
and the detectors trip at preset conductivity levels.  Last
replacement of the 10 IDZs was on 5/13/02 by Yuri Onufrienko, and one
of the units had been removed since then due to failure.]

CDR Kenneth Bowersox and FE-2/SO Don Pettit successfully conducted
the second Robotics operation this week.  [Their operations began at
about 10:45am with the "fast checkout" of the Tip LEE (latching end
effector) which could not be performed yesterday because the needed
SSRMS redundant string was unpowered due to power constraints.  They
then executed a ten-step single joint maneuver to move the arm into
position to grapple MBS PDGF1 (mobile base system power & data
grapple fixture #1), after which SSRMS switched its "base" from the
current Lab PDGF to MBS PDGF1 for subsequent performance of a series
of LEE operations to complete remaining LEE OCRs (on-orbit checkout
requirements).  As final task, the arm was maneuvered to the FMS
(force moment sensor) data gathering position #2. This placed theTip
LEE in a specific orientation which provides desired lighting
conditions to collect FMS data.  The latter were collected by the
ground at ops completion.] 

Nikolai Budarin worked on the Russian payload laptop 3, terminating
the cycling process of its two batteries, first one, later the
other. 

FE-2/SO Don Pettit activated the MSG (Microgravity Science Glovebox)
for a second day of operations with the InSPACE (Investigating the
Structure of Paramagnetic Aggregates from Colloidal Emulsions)
payload, today running tests 26 and 25.  [Purpose of InSPACE is to
obtain basic data on magnetorheological (MR) fluids, a new class of
"smart materials" or controllable fluids which can provide a quiet,
rapid-response interface between mechanical components and electronic
controls.  This makes them interesting for new brake systems, seat
suspensions, robotics, clutches, airplane landing gear and vibration
damping systems.  The experiment studies the formation of
three-dimensional microstructures from the small magnetic particles
in these fluids under the influence of pulsed magnetic fields inside
the experiment coil assemblies, in which the MR fluid is contained in
small precision rectangular borosilicate glass vials.]

VTR-2 (video tape recorder #2) is experiencing problems with
playback. VTR-1 was used for recording today,s InSPACE ops.  VTR-2
troubleshooting is underway.

CDR Kenneth Bowersox performed a lens change on the EarthKAM system
at the Lab window, going from 50mm to the 180mm-lens configuration.
[EarthKAM deactivation and stow is scheduled for 4/14 (Monday) and
re-setup and activation for 4/25 (Friday).  EK images can be viewed
on the Internet at
http://datasystem.earthkam.ucsd.edu/cgi-bin/datasys/ek_images_station

The Russian Flight Engineer performed more observation and
photography with the Russian GFI-8 Uragan ("hurricane") program, as
cloud cover permitted, us ing the Nikon D1 digital camera (f400 and
f800 lens).  Images from the CompactFlash card were later to be
downlinked via Regul packet mail.  [Areas of interest today were
cities in northern Iraq (to record environmental damages caused by
burning oil wells), the Kara Kum canal, long-range shots of the main
Karatau fault, Irtysh River flood plain, Central Altai mountains,
Kolka glacier north of Kazbek volcano, Chirkeisk reservoir, and the
Temirtau pits.] 

Pettit switched EMU batteries for capacity-preserving discharge,
terminating the process on the first and initiating it on the second
battery. 

Bowersox meanwhile worked a lengthy task on the cardiac defibrillator
system which required troubleshooting.  [Using a UOP (utility outlet
panel) power source and a voltage/current scopemeter, in part 1 of
the IFM (in-flight maintenance) he was to determine if the
defibrillator is leaking current from its battery when not in use.
Part 2 of the procedure was then to ascertain if the defibrillator's
battery charger is functioning properly.] 

Budarin terminated the bake-out/regeneration cycle for adsorption
bed #2 of the BMP micropurification unit and switched it back to
Purify mode.  Both BMP channels are now again running in this mode.
[The regeneration of the air purifier filter beds is repeated every
20 days.  Each bakeout to space vacuum takes about 24 hours.]

Afterwards, Nikolai removed two "Klest" TV cameras and their light
units from the descent module of the Soyuz TMA-1/5S. [The cameras are
not required for the return of the Expedition 6 crew, and their
removal increases 5S downmass (d/m) capability for science payloads.
US allocation will be about 4 kg, against a d/m request of 13 kg.]

Bowersox conducted another inventory audit of the available CWCs
(collapsible water containers) and their condition.

Sox also took care of the regular daily maintenance of the SOZh life
support system in the SM, while Don prepared the daily inventory
update file for downlink to the IMS (inventory management system)
database. 

The two U.S. crewmembers completed their weekly FFQs (food frequency
questionnaires) on  the MEC (medical equipment computer), which keep
a running account of their nutritional intake.

Budarin completed his regular daily 5-min. inspection of the BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

All crewmembers performed their regular daily physical exercise
program on RED, TVIS and VELO with load trainer.

At 4:40am EDT, the crew engaged in a live-TV downlink for a press
conference with Russian media assembled at TsUP for tomorrow's
Russian festive holiday (Cosmonautics Day, 4/12).  [Media represented
at the press conference were six television channels, three radio
stations, three news agencies, two trade magazines and two
newspapers.  Also present, on a time-permitting basis, were CNN
(USA), Times (UK), ARD (Germany) and Fuji (Japan).]

A conjunction with the TRMM satellite (Tropical Rainfall Measuring
Mission, object #25063) is predicted for three consecutive orbits
beginning 4/13 (Sunday) at 8:31pm EDT.  The TRMM has been encountered
before, last time on 2/21/03.  The conjunction is being monitored
closely to determine possible need for a DAM (debris avoidance
maneuver). 

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.